# Physics-Constrained Fine-Tuning of Flow-Matching Models for Generation and Inverse Problems

### v1

@@ 5 Conclusion
We have introduced a framework for post-training fine-tuning of flow-matching generative models to enforce physical constraints and jointly infer latent physical parameters informing the constraints. Through a novel architecture, combined with the combination of weak-form PDE residuals with an adjoint-matching scheme our method can produce samples that adhere to complex constraints without significantly affecting the sample diversity. Preliminary experiments on Darcy flow and acoustic wave problems demonstrate the potential of this method to reduce residuals and enable joint solution–parameter generation, supporting its promise for physics-aware generative modelling.
Future steps include adaptive approaches to optimising trade-off between constraint enforcement and generative diversity, and extending the framework to more complex and multi-physics systems, including coupled PDEs and stochastic or chaotic dynamics. We would also explore how this methodology can be leveraged for uncertainty quantification and propagation, and downstream tasks such as optimal sensor placement and scientific discovery workflows.
Acknowledgments
@@
This work was supported by funding from the German Federal Ministry for Education and Research (Bundesministerium für Bildung und Forschung, BMBF) within the project ’Physics Informed Anomaly Detection’ (grant number 01IS24071).
Appendix A Dataset Details
@@
In this section we detail the datasets used throughout our study. Our guiding principle was to select scenarios in which the underlying parameter fields contain sharp discontinuities, thereby inducing rich, non-linear behaviour in the associated state variables and making the inverse problem decidedly non-trivial. Although the Darcy‐flow benchmark follows the setup of Li et al. (2021), we regenerate the data so that the sample count, grid resolution, and ground-truth parameters can be controlled precisely. Complete scripts for producing both the Darcy and acoustic datasets will be released to facilitate transparency and reproducibility.
Darcy Flow Dataset.
@@
Each sample comprises a permeability field $a$ and its steady-state pressure solution $u$ on the unit square $\Omega=[0,1]^{2}$. The PDE
@@
$\nabla\!\cdot\!\bigl{(}a(\xi)\,\nabla u(\xi)\bigr{)}=f(\xi),\qquad$
is discretised on a $64\times 64$ grid with a five-point finite-difference stencil and harmonic averaging of $a$; the source is fixed to $f\equiv 1$. For the denoising experiments we consider $u|_{\partial\Omega}=0$, and for misspecified boundary conditions we set $u|_{\partial\Omega}(\xi_{1},1)=\sin(\pi\,\xi_{1})$.
Permeability field. A zero-mean Gaussian random field is generated by a Karhunen–Loève expansion with smoothness exponent $\alpha=2$ and correlation length $\tau=3$. To mimic sharply contrasting media, the field is mapped to a binary permeability: $a(\xi)=12$ if the raw GRF is non-negative and $a(\xi)=3$ otherwise.
Dataset Size. We compute 20.000 independent $(a,u)$ pairs, each stored at single-precision on the $64\times 64$ grid.
Samples from this dataset are shown in Fig. 5, Fig. 6, and Fig. 7 for the clean, noisy and misspecified setting respectively.
Figure 5: Samples from the (clean) Darcy dataset described in Appendix A.
Figure 6: Samples from the noisy Darcy dataset described in Appendix A.
Figure 7: Samples from the Darcy dataset using a sinusiodal boundary condition on the top edge described in Appendix A.
Acoustic Wave Dataset.
We model pressure propagation in a heterogeneous medium governed by
$\partial_{t}^{2}p(t,\xi)-c(\xi)^{2}\,\Delta p(t,\xi)$ $=0,\quad(\xi,t)\in\Omega\times[0,T]$
$\partial_{n}p(t,\xi)$ $=0,\quad(\xi,t)\in\partial\Omega\times[0,T]$
$p(0,\xi)$ $=h(\xi)$
with $\Omega=[0,1]^{2}$. The speed of sound $c(\xi)$ is obtained similarly to the permeability in the Darcy data set: a sample from a GRF is thresholded to produce a binary field. Here, the speed of sound takes the values $\{2,3\}$, and smoothness exponent and correlation length are set to $\alpha=2$ and $\tau=1$ respectively.
Initial Condition. The pressure at $t=0$ is a superposition of four identical isotropic Gaussians centred at $\bigl{(}\tfrac{i}{3},\tfrac{j}{3}\bigr{)}$ for $i,j\in\{1,2\}$:
$p(0,\xi)=\sum_{i,j=1}^{2}\!\exp\!\Bigl{(}-\frac{\|\xi-(i/3,j/3)\|^{2}}{2\sigma^{2}}\Bigr{)},$
with common variance $\sigma^{2}=10^{-2}$.
Numerical Solver. Spatial derivatives are discretised on a $64\times 64$ grid using a second-order centred Laplacian with reflective padding. Time integration employs the standard leap-frog scheme with step size $\Delta t=10^{-3}$. Solutions are recorded every ten five steps, yielding  64\,64 frames in total so that each sample forms a $64\times 64\times 64$ space–time tensor with a total time range of $t\in[0,\,0.315]$.
Dataset Size. We generate 10 000 independent $\bigl{(}p,c\bigr{)}$ pairs under the above settings, stored in single precision.
Figure 8: Two samples from the acoustic wave dataset described in Appendix A. The first column shows speed of sound samples from discretized GP which creates different patterns of wave propagation, starting from a superposition of four Gaussians. Evolution per row is shown for discrete time steps indexed by $i$.
Appendix B Pre-Training of Flow Matching Models
@@
We adopt the vanilla Flow–Matching (FM) procedure of Lipman et al. (2023). Based on the optimal-transport reference flow
@@
The mapping $v_{\theta}$ is realised with a U-FNO (Wen et al. 2022). Departing from the original design, we use the U-Net skip-connection structure in all layers. We prepend the physical input with fixed sinusoidal encodings for absolute spatial coordinates and the normalised time stamp so the purely spectral-convolutional backbone receives explicit space-time context without extra learnable parameters. Padding inside the U-FNO (for the spectral convolution layers) is reflective or replicate, matching the PDE’s boundary conditions. Before training, data is standardised to zero mean and unit variance.
@@
Pre-training is performed on a single NVIDIA L40S GPU (48GB GPU memory) with the hyperparameters listed in Tables 4 for 2D data and 5 for 3D data respectively. The optimiser is Adam with an initial learning rate of $3\times 10^{-4}$ (2D) or $5\times 10^{-4}$ (3D), halved at the epoch intervals specified in the tables. Training takes roughly six hours for the 2D model (500 epochs, batch size 256) and eight hours for the 3D model (50 epochs, batch size 8).
Table 4: U–FNO hyper-parameters used for 2D data.
Spectral modes / dim. 32
Hidden channels 32
Time emb. channels 32
Positional emb. channels 32
FNO layers 0
U–FNO layers 4
Batch size 256
Epochs 500
Initial learning rate $3\times 10^{-4}$
LR decay schedule $\times 0.5$ / 100 epochs
Parameters $\approx 17.4\text{M}$
Training time $\approx$ 6 h
Table 5: U–FNO hyper-parameters used for 3D (space–time) data.
Spectral modes / dim. 16
Hidden channels 32
Time emb. channels 30
Positional emb. channels 30
FNO layers 0
U–FNO layers 2
Batch size 8
Epochs 50
Initial learning rate $5\times 10^{-4}$
LR decay schedule $\times 0.5$ / 10 epochs
Parameters $\approx 34.6\text{M}$
Training time $\approx$ 8 h
Figures 9 and 10 present representative pressure–permeability pairs drawn from the Flow-Matching (FM) models trained on the noisy and misspecified Darcy datasets, respectively. Corresponding permeability fields were recovered with the pre-trained inverse predictor $\varphi$. Figure 11 shows analogous results for the acoustic‐wave dataset. In all three cases the generative models reproduce the salient structures of the data distribution, as evidenced by the fact that $\varphi$ can invert the generated states back to plausible parameter fields—a task it typically fails on when the synthetic samples deviate from the true manifold. The inverse predictions for the acoustic setting appear blurrier than those for Darcy. We attribute this to an architectural compromise made for computational efficiency, laid out in the next section: $\varphi$ inherently is designed as a 2D architecture that treats the 64 time frames as input channels, so temporal information is fused already in the lifting layer. A 3D backbone with later fusion stages would likely recover sharper time–independent sound-speed maps, albeit at higher computational cost.
Figure 9: Samples from a pre-trained FM model on the noisy Darcy dataset, including recovered permeabilities from a trained inverse predictor (right column).
Figure 10: Samples from a pre-trained FM model on the Darcy dataset with misspecified, including recovered permeabilities from a trained inverse predictor (right column).
Figure 11: Samples from a pre-trained FM model on the acoustic wave dataset, including recovered speed of sound coefficient fields from a trained inverse predictor. Temporal evolution is displayed for discrete time steps indexed by $i$.
Appendix C Method: Implementation Details

### v2

@@ 5 Conclusion
We have introduced a framework for post-training fine-tuning of flow-matching generative models to enforce physical constraints and jointly infer latent physical parameters informing the constraints. Through a novel architecture, combined with the combination of weak-form PDE residuals with an adjoint-matching scheme our method can produce samples that adhere to complex constraints without significantly affecting the sample diversity. Experiments across PDE problems demonstrate the potential of this method to reduce residuals and enable joint solution–parameter generation, supporting its promise for physics-aware generative modelling. Future steps include adaptive approaches to optimising trade-off between constraint enforcement and generative diversity, and extending the framework to more complex and multi-physics systems, including coupled PDEs and stochastic or chaotic dynamics. We would also explore how this methodology can be leveraged for uncertainty quantification and propagation, and downstream tasks such as optimal sensor placement and scientific discovery workflows.
Reproducibility Statement
We report datasets, model backbones, training schedules, loss definitions, evaluation metrics, and the key hyperparameters required to reproduce our results in the main text and appendix. Remaining implementation choices are documented in the released configuration files. Upon acceptance, we will open-source the code, including full training scripts. We fixed random seeds where applicable and specify hardware/software versions.
Acknowledgements
@@
This work was supported by funding from the German Federal Ministry for Education and Research (Bundesministerium für Bildung und Forschung, BMBF) within the project ’Physics Informed Anomaly Detection’ (grant number 16IS24071A).
Appendix A Use of Large Language Models
We employed large language models to polish the manuscript (wording, grammar, and synonyms) and to assist with implementation details, including plotting scripts and code rewriting/refactoring. The research questions, problem formulation, algorithmic design and experimental methodology, however, were conceived by the authors. All LLM-produced text and code were reviewed, adapted, and verified by the authors prior to inclusion.
Appendix B Dataset Details
@@
In this section we detail the datasets used throughout our study. Our guiding principle was to select scenarios in which the underlying parameter fields contain sharp discontinuities, thereby inducing rich, non-linear behaviour in the associated state variables and making the inverse problem decidedly non-trivial. Although the Darcy‐flow benchmark follows the setup of Li et al. (2020), we regenerate the data so that the sample count, grid resolution, and ground-truth parameters can be controlled precisely. Complete scripts for producing all data will be released to facilitate transparency and reproducibility.
B.1 Darcy Flow Dataset
@@
The dataset comprises 20,00020{,}000 independent samples, each a pair $(a,u)$ on the unit square $\Omega=[0,1]^{2}$, where $a:\Omega\to\mathbb{R}_{>0}$ is the permeability and $u:\Omega\to\mathbb{R}$ is the steady-state pressure solving
$\nabla\!\cdot\!\bigl(a(\xi)\,\nabla u(\xi)\bigr)=f(\xi),\qquad\xi=(\xi_{1},\xi_{2})\in\Omega,$
with homogeneous Dirichlet boundary conditions $u|_{\partial\Omega}=0$ and constant forcing $f\equiv 1$.
Discretization.
We use a uniform $64\times 64$ nodal grid with spacing $\Delta x=1/(64-1)$ and the standard five-point finite-difference scheme. Interface permeabilities are formed by two-point arithmetic averaging. Dirichlet values are imposed strongly, yielding a sparse SPD linear system that is solved with a direct sparse solver.
Permeability sampling.
We draw a zero-mean Gaussian random field $a_{\mathrm{raw}}$ via a cosine-basis Karhunen–Loève synthesis on $\Omega$ associated with the Matérn-type covariance operator
$\mathcal{C}\;=\;(\tau^{2}-\Delta)^{-\alpha},\qquad\alpha=2,\ \ \tau=3,$
i.e., using the DCT-II basis (Neumann eigenfunctions of $\Delta$), setting the DC mode to zero to enforce exact zero mean, scaling by the spectrum of $\mathcal{C}$, and applying an orthonormal inverse DCT to obtain a grid realization.
To model sharp contrasts, we threshold the Gaussian field into a piecewise-constant permeability,
$a(\xi)\;=\;\begin{cases}12,&a_{\mathrm{raw}}(\xi)\geq 0,\\[2.0pt] 3,&a_{\mathrm{raw}}(\xi)<0.\end{cases}$
Observational noise.
In experiments with noisy observations, we corrupt the pressure with additive Gaussian noise
$\tilde{u}\;=\;u+\sigma\,\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,I),\quad\sigma=10^{-4}.$
B.2 Linear Elasticity Dataset
The dataset contains N=10, [^0] N=10{,}000 independent samples on the unit square $\Omega=[0,1]^{2}$. Each sample is a pair $(E,u)$ where $E:\Omega\to\mathbb{R}_{>0}$ denotes the spatially varying Young’s modulus and $u:\Omega\to\mathbb{R}^{2}$ is the plane–strain displacement field under fixed Poisson ratio $\nu=0.30$. The boundary loading is deterministic and identical across samples: the left/right edges are clamped, and sinusoidal normal displacements are prescribed on the top and bottom edges with zero tangential slip (amplitudes $A_{\mathrm{top}}=A_{\mathrm{bot}}=0.10$). For each $E$, $u$ solves the static linear elasticity equations (no body force)
@@ B.2 Linear Elasticity Dataset
$\nabla\!\cdot\sigma(\xi)=0,\qquad\sigma\;=\;\lambda(E,\nu)\,\mathrm{tr}(\varepsilon)\,I+2\,\mu(E,\nu)\,\varepsilon,\qquad\varepsilon\;=\;\tfrac{1}{2}\bigl(\nabla u+\nabla u^{\!\top}\bigr),\quad\xi\in\Omega,$
with Lamé parameters $\lambda(E,\nu)=\tfrac{\nu\,E}{(1+\nu)(1-2\nu)}$ and $\mu(E,\nu)=\tfrac{E}{2(1+\nu)}$, and Dirichlet boundary conditions
$u(\xi_{1}{=}0,\xi_{2})=0,\quad u(\xi_{1}{=}1,\xi_{2})=0$
$u_{x}(\xi_{1},\xi_{2}{=}0)=0,\ \ u_{y}(\xi_{1},\xi_{2}{=}0)=-A_{\mathrm{bot}}\sin(\pi\,\xi_{1})$
$u_{x}(\xi_{1},\xi_{2}{=}1)=0,\ \ u_{y}(\xi_{1},\xi_{2}{=}1)=+A_{\mathrm{top}}\sin(\pi\,\xi_{1}).$
During fine-tuning, we set $A_{\mathrm{bot}}=0.075$ to enforce adaptation of the solutions, simulating a misspecification between observed data and the assumed model.
Discretization.
We use a uniform $64\times 64$ grid on $\Omega$ and standard second–order finite differences in the small–strain regime. Dirichlet data are imposed strongly. The discrete equilibrium equations are advanced by a stable explicit time–marching (damped gradient) scheme for the static problem until the global residual norm falls below $10^{-6}$, or a cap of $2\times 10^{4}$ iterations is reached.
Coefficient field sampling (piecewise-constant Voronoi medium).
Heterogeneous modulus fields $E$ are obtained from a Voronoi tessellation constructed by drawing a fixed number of sites uniformly in $\Omega$, assigning to each Voronoi cell a modulus sampled log-uniformly within [1.0, 10.0][1.0,\,10.0], and rasterizing the resulting partition to the computational grid by nearest-site labeling. This produces piecewise-constant $E$ with sharp jumps that emulate multi-phase media. To control interface smoothness, a separable Gaussian blur with standard deviation $\sigma_{\mathrm{blur}}=1.0$ (in grid units) is applied to the rasterized field.
B.3 Helmholtz Wave Propagation Dataset
The dataset consists of N=10, [^0] N=10{,}000 independent samples on the unit square $\Omega=[0,1]^{2}$. Each sample is a pair $(c,u)$, where $c:\Omega\to\mathbb{R}_{>0}$ denotes the spatially varying sound speed and $u:\Omega\to\mathbb{C}$ is the time–harmonic acoustic pressure. For a fixed angular frequency $\omega>0$, the field $u$ solves the damped Helmholtz equation
$\Delta u(\xi)-\bigl(1-i\tan\delta\bigr)\,\kappa(\xi)^{2}\,u(\xi)\;=\;s(\xi),\qquad\kappa(\xi)\;=\;\frac{\omega}{c(\xi)},\qquad\xi\in\Omega,$
subject to homogeneous Robin boundary conditions
$\partial_{n}u(\xi)+i\,\gamma\,u(\xi)=0,\qquad\xi\in\partial\Omega,$
with fixed admittance parameter $\gamma>0$ and loss tangent $\tan\delta\geq 0$.
Discretization.
We use a uniform $64\times 64$ nodal grid on $\Omega$ with spacing $\Delta x=1/(64-1)$. The Laplacian $\Delta$ is discretized by the standard five-point finite-difference stencil in the interior and homogeneous Neumann conditions in the normal direction on all sides, implemented via ghost-point elimination. The Robin condition $\partial_{n}u+i\gamma u=0$ is enforced by adding purely imaginary diagonal contributions at boundary nodes to the stiffness matrix, consistent with the underlying Neumann-type discretization. The heterogeneous reaction term $(1-i\tan\delta)\kappa^{2}u$ is assembled pointwise on the grid. This yields a sparse complex-valued linear system, which is solved for each sample using a direct sparse solver.
Sound-speed sampling.
The sound speed $c$ is derived from a Gaussian random field constructed in the same manner as for the Darcy dataset, but interpreted as a wave speed. We first draw a zero-mean Gaussian field $g$ on the grid using a cosine-basis Karhunen–Loève synthesis associated with the Matérn-type covariance operator
$\mathcal{C}\;=\;(\tau^{2}-\Delta)^{-\alpha},\qquad\alpha=4,\ \ \tau=6,$
using the DCT-II basis on $\Omega$, zeroing the DC mode, scaling by the spectrum of $\mathcal{C}$, and applying the inverse DCT to obtain $g$. To model sharp material contrasts, we map $g$ to a two-level sound speed,
$c(\xi)\;=\;\begin{cases}c_{\max},&g(\xi)\geq 0,\\[2.0pt] c_{\min},&g(\xi)<0,\end{cases}$
with $c_{\min}=0.8$ and $c_{\max}=1.2$ in the reported experiments.
Source and damping.
The source term $s$ is a fixed, smooth Gaussian bump centered in $\Omega$,
$s(\xi)\;=\;\exp\!\Bigl(-\tfrac{\|\xi-(0.5,\,0.5)\|_{2}^{2}}{2\sigma^{2}}\Bigr),\quad\sigma=0.05,$
shared across all samples. We use a fixed angular frequency $\omega=20$ and Robin parameter $\gamma=0.03$. Data is generated with a small but nonzero loss tangent $\tan\delta=0.02$, leading to complex-valued damped solutions.
B.4 Stokes Lid-Driven Cavity Dataset
The dataset consists of N=10, [^0] N=10{,}000 independent samples on the unit square $\Omega=[0,1]^{2}$. Each sample is a pair $(\nu,u)$, where $\nu:\Omega\to\mathbb{R}_{>0}$ denotes the spatially varying kinematic viscosity and $u:\Omega\to\mathbb{R}^{2}$ is the steady incompressible velocity field with associated pressure $p:\Omega\to\mathbb{R}$. For each $\nu$, $(u,p)$ solve the variable-coefficient Stokes system
[…]
with a smooth lid-driven cavity boundary condition on the top wall
$u(\xi_{1},\xi_{2}{=}1)\;=\;\bigl(U_{\mathrm{lid}}\sin^{2}(\pi\,\xi_{1}),\,0\bigr),$
and no-slip conditions on the remaining walls
$u(\xi_{1},\xi_{2}{=}0)=0,\qquad u(\xi_{1}{=}0,\xi_{2})=0,\qquad u(\xi_{1}{=}1,\xi_{2})=0.$
To excite richer flow structures, we add a Kolmogorov-type body force
$f(\xi_{1},\xi_{2})\;=\;F_{0}\bigl(\sin(\pi\,\xi_{2}),\,0\bigr),$
with fixed amplitude $F_{0}>0$.
Discretization.
The Stokes system is discretized using FEniCS (Baratta et al., 2023; Scroggs et al., 2022b; a; Alnaes et al., 2014) on a conforming triangular mesh obtained from a $(64{-}1)\times(64{-}1)$ grid of squares, each split into two triangles. We employ Taylor–Hood elements (vector-valued $P_{2}$ for $u$, scalar $P_{1}$ for $p$) on the mixed space $V\times Q$. The weak form reads
$\int_{\Omega}\nu\,\nabla u:\nabla v\,dx-\int_{\Omega}p\,\nabla\!\cdot v\,dx-\int_{\Omega}q\,\nabla\!\cdot u\,dx\;=\;\int_{\Omega}f\cdot v\,dx,$
for all test functions $(v,q)\in V\times Q$, with the lid and no-slip boundary conditions imposed strongly. For each sample, the resulting saddle-point system is solved with a direct solver, and the solution $(u,p)$ is sampled on a uniform $64\times 64$ grid on $\Omega$; the pressure is post-processed to have zero spatial mean in order to fix the gauge.
Viscosity field sampling.
The viscosity $\nu$ is obtained from a Gaussian random field constructed in the same manner as for the Darcy and Helmholtz datasets. We first draw a zero-mean Gaussian field $g$ on the grid using a cosine-basis Karhunen–Loève synthesis associated with the Matérn-type covariance operator
$\mathcal{C}\;=\;(\tau^{2}-\Delta)^{-\alpha},\qquad\alpha=4,\ \ \tau=6,$
using the DCT-II basis on $\Omega$, zeroing the DC mode, scaling by the spectrum of $\mathcal{C}$, and applying the inverse DCT to obtain $g$. To model sharp viscosity contrasts, we map $g$ to a two-level field,
$\nu(\xi)\;=\;\begin{cases}\nu_{\max},&g(\xi)\geq 0,\\[2.0pt] \nu_{\min},&g(\xi)<0,\end{cases}$
with $\nu_{\min}=0.02$ and $\nu_{\max}=0.5$ in the reported experiments.
Lid and body-force parameters.
In all samples we fix the lid velocity scale to $U_{\mathrm{lid}}=1.0$ and the Kolmogorov forcing amplitude to $F_{0}=2.0$.
Appendix C Pre-Training of Flow Matching Models
@@
We adopt the vanilla Flow–Matching (FM) procedure of Lipman et al. (2023). Based on the reference flow
@@
For PDE data, the mapping $v_{\theta}$ is realised with a U-FNO (Wen et al., 2022), which combines Fourier Neural Operator (FNO, Li et al. (2020)) with U–Net (Ronneberger et al., 2015) layers. The FNO layers have an inductive bias towards low-frequency solutions and are therefore particularly suited for modeling PDE data, while the U–Net layers help to refine outputs and produce discontinuities.
Departing from the original design, we use the U–Net skip-connection structure in all layers. Also, we employ a more powerful time-conditioning in the U–Net layers by using FiLM-style Adaptive Group Normalization (AdaGN) time conditioning, i.e., predicting per-channel scale and shift from a sinusoidal time embedding and applying them after GroupNorm in each residual block. This follows the feature-wise linear modulation idea of Perez et al. (2018) and its diffusion U–Net instantiations in Dhariwal and Nichol (2021); Rombach et al. (2022).
We prepend the physical input with fixed sinusoidal encodings for absolute spatial coordinates and the normalised time stamp so the spectral backbone receives explicit space-time context. Padding inside the U-FNO (for the spectral convolution layers) is reflective, which has empirically worked better than replicative. Before training, data is standardised to zero mean and unit variance.
Table 3 compiles the hyperparameters of the U-FNO architecture used for the base FM models in our experiments.
Table 3: U–FNO (2D) backbone hyperparameters used in our experiments.
Hyperparameter Value Description
Number of layers 4 Count of spectral and U–Net operator blocks.
Padding mode reflect Boundary padding for convolutions/lifts.
Input channels 1 - 3 Channels of input field $x$.
Output channels 1 - 3 Channels of output field.
Time embedding dim 32 Dimensionality of time conditioning.
Spatial embedding dim 32 Dimensionality of coordinate embedding.
Lifting width 256 Channels in input lifting/projection.
Fourier modes (kx, ky) [32, 32] Retained spectral modes per axis.
Spectral block width 32 Channel width within FNO layers.
U–Net base widths 32 Stage-wise channel widths.
Embedding width (U–Net) 64 Channels for conditioning embeddings.
Attention stages [] Stages with self-attention (empty $\Rightarrow$ none).
Attention heads [] Multi-head count if attention is enabled.
Total number of parameters $\approx$ 19M
On natural image data, we use a pretrained Latent Flow Matching Model (Dao et al., 2023) based on a DiT (Peebles and Xie, 2023) backbone.
@@
We train the FM backbone with AdamW (Loshchilov and Hutter, 2017), using a linear warmup of the learning rate from 0 to the base value over the first $p_{\mathrm{wu}}$ fraction of training steps, followed by a constant learning rate thereafter. For evaluation stability, we maintain an exponential moving average (EMA) of the network parameters $\theta^{\mathrm{EMA}}\leftarrow\beta\,\theta^{\mathrm{EMA}}+(1-\beta)\,\theta$, a practice rooted in Polyak averaging (Polyak and Juditsky, 1992) and widely adopted in modern deep models (e.g., Tarvainen and Valpola, 2017).
Table 4: Flow Matching (FM) training hyperparameters and schedule.
Hyperparameter Description
Batch size 128 Minibatch size per step.
Base learning rate 1e-4 AdamW step size.
Warmup fraction $p_{\mathrm{wu}}$ 0.01 Fraction of total steps used for linear LR warmup.
Epochs 300 Full passes over the dataset.
Optimizer AdamW Applied to FM backbone parameters.
Weight decay 0.01 AdamW L2/decoupled decay coefficient.
EMA decay 0.9998 Exponential moving average of weights for evaluation.
Schedule after warmup constant LR held constant after warmup.
The FM training configuration is summarized in Table 4. For the Darcy dataset (20k samples), training takes around 12 hours on a single NVIDIA RTXA6000 GPU. For the smaller datasets of 10k samples but with more channels (Elasticity, Helmholtz, Stokes), training takes around 7 hours using the same configuration.
Appendix D Method: Implementation Details